# Intramolecular Singlet Fission Through a Coherently Coupled Excimer-like Intermediate

### v1

@@ Abstract
Singlet Fission (SF) into two triplets offers exciting avenues for high-efficiency photovoltaics and optically initializable qubits. While the chemical space of SF chromophores is ever-expanding, the mechanistic details of electronic-nuclear motions that dictate the photophysics are unclear. Rigid SF dimers with well-defined orientations are necessary to decipher such details. Here, using polarization-controlled white-light two-dimensional and pump-probe spectroscopies, we investigate a new class of contorted naphthalenediimide dimers, recently reported to have a favorable intramolecular SF (iSF) pathway. 2D cross-peaks between the two Davydov components and strongly wavelength-dependent $TT_{1}$ formation kinetics together implicate a coherently coupled excimer-like intermediate. Enhanced quantum beats in the $TT_{1}$ photoproduct suggest that inter-chromophore twisting and ruffling motions drive the $\sim$200 fs evolution towards excimer-like intermediate and the subsequent $\sim$2 ps relaxation to the $TT_{1}$ photoproduct. Polarization anisotropy directly tracks electronic motion during these steps and reveals surprisingly minimal electronic reorientation with significant singlet-triplet mixing throughout the nuclear evolution away from the Franck-Condon geometry towards relaxed $TT_{1}$. The combination of white-light 2DES and polarization anisotropy experiments presented here lends powerful insights for refinement of the existing theoretical models for iSF, and lays the foundation for a broader mechanistic investigation of iSF across various molecular platforms.
@@ 1 Introduction
Structurally rigid intramolecular SF (iSF) dimers are a particularly interesting testbed for understanding the mechanistic aspects of SF because exquisite control over electronic coupling matrix elements becomes possible by synthetically tuning molecular symmetry17. One of the pitfalls of structurally well-defined dimers is that excessive orbital overlaps due to center-to-center distances of only 3–4 Å can also lead to the undesired formation of excimers18 and low-energy $CT$ traps19. Consequently, the interplay of molecular geometry, electronic couplings and possibly even nuclear motions that differentiates one photophysical outcome from another becomes quite challenging to decipher. For example, $CT$ states are reported to both, mediate10, 11 and sometimes inhibit19 SF. In the context of iSF, excimer formation has been exclusively reported as an undesired pathway11, 18, 20 that is enabled by structural relaxation11, 21 on the excited state leading to a frustrated photodimerization which increases22 with electronic coupling. When SF is intermolecular, excimer-mediated SF is rare and is reported to be quite slow – 2.2 ps excimer formation and 22 ps $TT_{1}$ formation in diketopyrrolopyrrole thin films23, 46 ps $TT_{1}$ formation in $\pi$-stacked terrylenediimide (TDI) thin films10, and 2–7 ps $TT_{1}$ formation in TIPS-Tc nanoaggregates24 where in the latter two cases the excimer formation was faster than the few hundred femtosecond instrument response.
@@
Theoretical models from Krylov21 and co-workers have accounted for excimer formation to argue that structural relaxation of the photoexcited singlets in flexible staggered tetracene dimers towards an excimer minima can in fact make the $TT_{1}$ adiabat energetically accessible compared to the Franck-Condon (FC) geometry, but often at the expense of inhibiting the $TT_{1}$ formation by enhancing non-radiative relaxation of the excimer. The latter parasitic pathway is in line with the observations11 of Hong et al. who report that a structurally relaxed iSF dimer (excimer, $EX$) can borrow $S_{1}$ oscillator strength and relax directly to the ground state. Schatz and co-workers have extended the $CT$-mediated picture of SF to propose12 a state vector picture for various configurations that may be present in iSF dimers such that the excimer $EX$ is defined as having a borrowed electronic character from $LE$ and $CT$ states and possibly also from the $TT_{1}$ state depending on their relative adiabatic energies. At the same time, recent observations of near-instantaneous generation of coherently mixed LECTLE-CT intermediates25 in perylenediimide (PDI) trimers, vibronically coherent formation of LETTLE-$T$ intermediates26 during intermolecular SF in tetracene thin films, and EXCTEX-CT non-adiabatic intermediates27 during symmetry breaking charge separation (SBCS) dimers prompted us to explore the following questions in the context of iSF – 1. Can coherent intermediates, formed as a result of strong electronic/vibronic couplings, overcome the unavoidable excimer traps to form $TT_{1}$ in strongly coupled iSF dimers? 2. Davydov splittings and therefore wavelength dependent iSF is theoretically expected16 in strongly coupled iSF dimers but both features have not been reported in any iSF system till date. 3. What is the electronic reorientation involved during the evolution of a photoexcited singlet towards these intermediates? Given the near-instantaneous generation of above intermediates, this question remains unclear and has not yet been probed directly.
@@
In relation to the latter question, our recent work16 on a vibronic dimer picture of iSF with explicit role of low and high-frequency vibrations reveals that given the large nuclear reorganization energies on the $TT_{1}$ state, only a small amount of electronic mixing between $LE$ and $TT_{1}$ – direct or mediated – is sufficient to lead to vibronically enhanced $LE-TT_{1}$ mixing. Given this expectation, rich dynamical effects such as coherent generation of a $LE-TT_{1}$ intermediate and wavelength dependent iSF rates depending on the excitation of lower or upper Davydov component of the dimer, are to be expected but not reported for any iSF system till date. One of us28 recently reported a new class of naphthalenediimide (NDI) dimers where a contortion of the NDI skeleton along with a relative rotation of 20 and 24 along the long and short axes, respectively, avoids excimer formation even with center-to-center distance of 3.3 Åbetween the rings, such that iSF becomes a favorable pathway. Sub-picosecond iSF was confirmed by narrowband low-temporal resolution pump-probe (PP) experiments and through a Ruthenium-based triplet sensitizer with NDI dimer triplets directly probed by microsecond PP experiments. Absence of any $CT$ intermediates was confirmed by absence of radical anion and cation NDI bands29 in the PP spectra. The questions in the preceding paragraph are best explored in rigid strongly coupled iSF dimers and prompted us for a deeper mechanistic investigation of iSF in this system.
@@
Here, using a combination of white-light polarization-controlled two-dimensional electronic spectroscopy (2DES) and impulsive PP experiments we show that iSF in NDI dimers proceeds through an excimer-like intermediate that is coherently coupled to the initially excited singlet and forms on a $\sim$200 fs timescale. This coherent picture of iSF is evidenced by our measurements 2DES cross-peaks (CPs) and the reduced electronic polarization anisotropy, where both observations are consistent with strong electronic correlations between the chromophore sites. Interestingly, the formation of this intermediate is not excitation wavelength dependent. However, its subsequent relaxation to the $TT_{1}$ state is strongly wavelength dependent – 2 ps when upper Davydov component is excited versus 0.6 ps when the lower Davydov component is excited. Enhanced quantum beats observed in the $TT_{1}$ photoproduct suggests that inter-chromophore twisting and ruffling motions likely drive the excited state evolution towards the excimer-like intermediate and the subsequent $TT_{1}$ formation. By directly tracking the electronic reorientation during $S_{1}$, $S_{1}+TT_{1}]$ intermediate and $TT_{1}$ formation, we find that that there is surprisingly little electronic reorientation during these steps and that strong electronic correlations between the chromophore sites and significant singlet-triplet mixing are maintained throughout the $TT_{1}$ formation process. Together, our observations represent the first instance where a strongly wavelength-dependent coherent formation of an excimer-like intermediate and subsequent $TT_{1}$ formation is reported in iSF. Our observations also demonstrate that tracking electronic transition dipole reorientation during $TT_{1}$ formation can provide powerful insights to improve the existing theoretical models for SF, and lay the foundation for a broader mechanistic investigation of iSF across various molecular platforms.
@@ 2.1 Linear Measurements
2.2 2DES reveals Davydov splittings and excitation-wavelength dependent $TT_{1}$ formation
@@
A 2DES map correlates excitation and detection frequencies of a system in the form of a 2D contour map of detection versus excitation axis which evolves as a function of pump-probe waiting time $T$. Our 2DES experiments are conducted in a partially collinear pump-probe geometry where the pump-pulse pair is generated through a common path interferometer. The collected signal measures the absorptive changes in the transmitted probe spectrum. Further experimental details are provided in Section S1. To investigate the nature of the absorption peaks at 600 nm and 650 nm ( Figure 1a), we conducted 2DES measurements of kinetic rate maps (RMs). Figure 2a shows the NDI dimer spectrum at $T$ = 0.2 ps. A clear cross-peak (CP) with excitation wavelength 650 nm and detection wavelength 608 nm, that is, between the main absorption peak and its red shoulder is seen. This is denoted by the white square in Figure 2a.
@@
Figure 2: 2DES reveals Davydov splittings and excitation wavelength dependent $TT_{1}$ formation. (a) 2D spectrum of the NDI dimer at $T$ = 0.2 ps. The top panel shows the absorption spectrum (black) overlaid with the pump spectrum (gray shaded). The side panel shows the spectrally integrated spectrum versus the detection wavelength, $\lambda_{t}$, overlaid with the probe spectrum (gray shaded). Contours are drawn at 10% to 100% in 10% intervals. The positive ground state bleach (GSB) signal is shown in red while the negative excited state absorption (ESA) signal is shown in blue. (b) (left) Dynamics in the ESA region denoted by the green rectangle in panel A. (middle) Dynamics in the ESA region denoted by the yellow rectangle in panel A. (right) Dynamics in the ESA region denoted by dashed yellow rectangle in panel A. For panel b (left and middle), the data was collected until 5 ps and the spectrally integrated 2D signals are fitted to a freely floated 3-exponential model where the first two time-constants are shown in the figure. For panel b (right), the time constants of the fit function for the middle panel are kept fixed and only the amplitudes are floated. The accompanying top panel shows the respective residuals. The $\pm{\sigma}$ error band on the fit is calculated by averaging N=6 trials of $T$ = 0.5 ps 2D spectrum where collection of individual trials was interleaved over the duration of the full scan. The arrows denote the 200 fs decay or rise, that is followed by the slower rise of the ESA signal. (c) 2D rate maps obtained by global fitting of all 2D pixels to a 3-exponential model. The obtained time constants are shown above the respective maps. The rate maps are normalized by the sum of the maxima of two rate maps. Global rate analysis neglects the excitation wavelength dependence seen in panel b (left versus middle). Contours are drawn from 10% to 100% in 10% intervals.
It is important to note that this is a positive upper 2D cross-peak ($\lambda_{t}<\lambda_{\tau}$, CPU) which arises only for strongly correlated electronic transitions where bleaching one transition also affects the other, that is, a shared ground state between electronic states. Thus, the presence of a positive CPU is a direct evidence for the main peak and the red shoulder in the absorption spectrum being Davydov components with a $\sim$1040 cm-1 energetic splitting suggesting large electronic coupling between the sites. Unlike conformationally heterogeneous18, 30 iSF dimers, this Davydov splitting is easily discernible in the structurally well-defined rigid dimer investigated here.
@@
Interestingly, the strongest features in the 2D spectrum are two the distinct ESA bands along the excitation axis, where both approximately coincide with the position of the two Davydov components. As shown in Figure 2b(left, middle), when an area at these locations is sampled, green versus yellow rectangles in Figure 2a, a strong wavelength-dependent dynamics of the ESA signal is seen. The ESA signal shows a fast decay of 0.19 $\pm$ 0.05 ps and 0.17 $\pm$ 0.05 ps with green versus yellow excitation, which is approximately wavelength-independent within the error bars. The rise of the ESA signal is, however, $\sim$3x faster when the lower Davydov component is excited (2.12 $\pm$ 0.32 ps versus 0.60 $\pm$ 0.18 ps). The fitting procedure is described in Section S2 and summarized in Table S1, S2. The fits shown in Figure 2b (left, middle) are freely floated and do not show any trends in the residual (top panels in Figure 2b). In comparison, when the fit is constrained to have the same time constants, a clear residual trend is seen in the 650 nm excitation wavelength region. This is shown in Figure S3b(right). Figure 2b(right) shows the $T$ dynamics in the dashed red rectangle with excitation wavelength at the lower Davydov component as before, but with detection wavelength $\sim$710 nm compared to 760 nm earlier. The $\sim$200 fs decay changes to a $\sim$200 fs rise of the signal, followed by a slower rise on the $\sim$2 ps timescale. Previous28 low temporal resolution PP experiments identified the picosecond rise of the ESA signal as the signature of $TT_{1}$ formation. Here, we find that this process is strongly wavelength dependent, along with an intermediate step that occurs on a $\sim$200 fs timescale, which, however, is not wavelength dependent. Note that we have confirmed that all the above 2DES observations remain unchanged even if the experiments were conducted in a non-polar (weakly polar) DCM solvent ( Section S5) suggesting that the intermediate states during $S_{1}\rightarrow TT_{1}$ do not involve any significant $CT$ character (vide infra).

### v3

@@ Abstract
Singlet Fission (SF) into two triplets offers exciting avenues for high-efficiency photovoltaics and optically initializable qubits. While the chemical space of SF chromophores is ever-expanding, the mechanistic details of electronic-nuclear motions that dictate the photophysics are unclear. Rigid SF dimers with well-defined orientations are necessary to decipher such details. Here, using polarization-controlled white-light two-dimensional and pump-probe spectroscopies, we investigate a new class of contorted naphthalenediimide dimers, recently reported to have a favorable intramolecular SF (iSF) pathway. 2D cross-peaks directly identify the two Davydov components of the dimer along with strongly wavelength-dependent $TT_{1}$ formation kinetics depending on which Davydov component is excited, implicating a coherently coupled intermediate that mediates iSF. Enhanced quantum beats in the $TT_{1}$ photoproduct suggest that inter-chromophore twisting and ruffling motions drive the $\sim$200 fs evolution towards an excimer-like intermediate and its subsequent $\sim$2 ps relaxation to the $TT_{1}$ photoproduct. Polarization anisotropy directly tracks electronic motion during these steps and reveals surprisingly minimal electronic reorientation with significant singlet-triplet mixing throughout the nuclear evolution away from the Franck-Condon geometry towards relaxed $TT_{1}$. The observations of coherent excimer-like intermediate and significant singlet-triplet mixing throughout the iSF process need to be carefully accounted for in the synthetic design and electronic structure models for iSF dimers aiming for long-lived high-spin correlated triplets.
@@ 1 INTRODUCTION
Structurally rigid intramolecular SF (iSF) dimers are a particularly interesting test bed for understanding the mechanistic aspects of SF because exquisite control over electronic coupling matrix elements becomes possible by synthetically tuning molecular symmetry17. One of the pitfalls of structurally well-defined dimers is that excessive orbital overlaps due to center-to-center distances of only 3–4 Å can also lead to the undesired formation of excimers18 and low-energy $CT$ traps19. Consequently, the interplay of molecular geometry, electronic couplings and possibly even nuclear motions that differentiates one photophysical outcome from another becomes quite challenging to decipher. For example, $CT$ states are reported to both, mediate10, 11 and sometimes inhibit19 SF. In the context of iSF, excimer formation has been generally reported as an undesired pathway11, 18, 20, 21 that is fast enough to effectively compete against the slower $TT_{1}$ formation kinetics. Excimer formation is reported to be enabled by structural relaxation11, 22, 23, 24 on the excited state leading to a frustrated photodimerization which increases25 with electronic coupling. Only fairly recent works report26, 21 excimer-mediated iSF but with slow $TT_{1}$ formation on $\sim$100 ps and slower timescales. When SF is intermolecular, excimer-mediated SF is reported to be similarly slower – 2.2 ps excimer formation and 22 ps $TT_{1}$ formation in diketopyrrolopyrrole thin films27, 46 ps $TT_{1}$ formation in $\pi$-stacked terrylenediimide (TDI) thin films10, and 2–7 ps $TT_{1}$ formation in TIPS-Tc nanoaggregates28 where in the latter two cases the excimer formation was faster than the few hundred femtosecond instrument response. Kim, Würthner and co-workers have elucidated23, 24, 21 the structural dynamics associated with the photophysics of excimer formation. The corresponding electronic dynamics carries similar mechanistic significance and is one of the focus of this work.
@@
Theoretical models from Krylov22 and co-workers have accounted for excimer formation to argue that structural relaxation of the photoexcited singlets in flexible staggered tetracene dimers towards an excimer minima can in fact make the $TT_{1}$ adiabat energetically accessible compared to the Franck-Condon (FC) geometry, but often at the expense of inhibiting the $TT_{1}$ formation by enhancing non-radiative relaxation of the excimer. The latter parasitic pathway is in line with the observations11 of Hong et al. who report that a structurally relaxed iSF dimer (excimer, $EX$) can borrow $S_{1}$ oscillator strength and relax directly to the ground state. Schatz and co-workers have extended the $CT$-mediated picture of SF to propose12 a state vector picture for various configurations that may be present in iSF dimers such that the excimer $EX$ is defined as having a borrowed electronic character from $LE$ and $CT$ states and possibly also from the $TT_{1}$ state depending on their relative adiabatic energies. At the same time, recent observations of near-instantaneous generation of coherently mixed LECTLE-CT intermediates29 in perylenediimide (PDI) trimers, vibronically coherent formation of LETTLE-$T$ intermediates30 during intermolecular SF in tetracene thin films, and EXCTEX-CT non-adiabatic intermediates31 during symmetry breaking charge separation (SBCS) dimers prompted us to explore the following questions in the context of iSF – 1. Can coherent intermediates, formed as a result of strong electronic/vibronic couplings, overcome the unavoidable excimer traps to form $TT_{1}$ in strongly coupled iSF dimers? 2. Strongly coupled sites and wavelength dependent iSF is theoretically expected16 in rigid iSF dimers but both features have not been directly confirmed in any iSF system till date. 3. What is the electronic reorientation involved during the evolution of a photoexcited singlet towards these intermediates? Given the near-instantaneous generation of above intermediates, this question remains unclear and has not yet been probed directly.
@@
In relation to the latter question, our recent work16 on a vibronic dimer picture of iSF with explicit role of low and high-frequency vibrations reveals that given the large nuclear reorganization energies on the $TT_{1}$ state, only a small amount of electronic mixing between $LE$ and $TT_{1}$ – direct or mediated – is sufficient to lead to vibronically enhanced $LE-TT_{1}$ mixing. Given this expectation, rich dynamical effects such as coherent generation of a $LE-TT_{1}$ intermediate and wavelength dependent iSF rates depending on the excitation of lower or upper Davydov component of the dimer, are to be expected but not reported for any iSF system till date. One of us32 recently reported a new class of naphthalenediimide (NDI) dimers where a contortion of the NDI skeleton along with a relative rotation of 20 and 24 along the long and short axes, respectively, avoids excimer formation even with center-to-center distance of 3.3 Åbetween the rings, such that iSF becomes a favorable pathway. Sub-picosecond iSF was confirmed by non-impulsive pump-probe (PP) experiments and through a Ruthenium-based triplet sensitizer with NDI dimer triplets directly probed by microsecond PP experiments. Absence of any $CT$ intermediates was confirmed by absence of radical anion and cation NDI bands33 in the PP spectra. The questions in the preceding paragraph are best explored in rigid strongly coupled iSF dimers and prompted us for a deeper mechanistic investigation of iSF in this system.
@@
Here, using a combination of white-light, polarization-controlled two-dimensional electronic spectroscopy (2DES) and impulsive PP experiments we show that iSF in NDI dimers proceeds through a coherently coupled intermediate that forms on a $\sim$200 fs timescale. This coherent picture of iSF is evidenced by our measurements of 2DES cross-peaks (CPs) and the reduced electronic polarization anisotropy, where both observations are consistent with strong electronic correlations between the chromophore sites. Interestingly, the formation and rate of this $S_{1}+TT_{1}]$ intermediate is not excitation wavelength or solvent polarity dependent. However, its subsequent relaxation to the $TT_{1}$ state is strongly wavelength dependent – 2 ps when upper Davydov component is excited versus 0.6 ps when the lower Davydov component is excited. Enhanced quantum beats observed in the $TT_{1}$ photoproduct suggest that inter-chromophore twisting and ruffling motions likely drive the excited state evolution towards an excimer-like intermediate and subsequent to a relaxed $TT_{1}$ state. By directly tracking the electronic reorientation during $S_{1}$, $S_{1}+TT_{1}]$ intermediate and $TT_{1}$ formation, we find that there is surprisingly little electronic reorientation during these steps and that strong electronic correlations between the chromophore sites and significant singlet-triplet mixing are maintained throughout the $TT_{1}$ formation process. To best of our knowledge, the observations of 2DES cross-peaks, coherent excimer-like intermediate, and strongly wavelength-dependent iSF with surprisingly minimal electronic reorientation throughout the process have not been reported before in any iSF system and elucidate the deeper mechanistic landscape that underpins singlet fission. Our observations also demonstrate that tracking electronic transition dipole reorientation during $TT_{1}$ formation can provide powerful insights to advance the synthetic design and electronic structure models of strongly coupled iSF dimers aiming to synthetically tune long-lived high-spin triplet states.
@@ 2.1 Linear Measurements
2.2 2DES reveals Davydov splitting and excitation-wavelength dependent $TT_{1}$ formation
@@
A 2DES map correlates excitation and detection frequencies of a system in the form of a 2D contour map of detection versus excitation axis which evolves as a function of pump-probe waiting time $T$. Our 2DES experiments are conducted in a partially collinear pump-probe geometry where the pump-pulse pair is generated through a common path interferometer. The collected signal measures the absorptive changes in the transmitted probe spectrum. Further experimental details are provided in Section S1. To investigate the nature of the absorption peaks at 600 nm and 650 nm ( Figure 1a), we conducted 2DES measurements of kinetic rate maps (RMs). Figure 2a shows the NDI dimer spectrum at $T$ = 0.2 ps. A clear cross-peak (CP) with excitation wavelength 650 nm and detection wavelength 608 nm, that is, between the main absorption peak and its red shoulder is seen. This is denoted by the white square in Figure 2a. It is important to note that this is a positive upper 2D cross-peak ($\lambda_{t}<\lambda_{\tau}$, CPU) which arises only for strongly correlated electronic transitions where bleaching one transition also affects the other, that is, a shared ground state between electronic states. Thus, the presence of a positive CPU is a direct evidence for the main peak and the red shoulder in the absorption spectrum being Davydov components with a $\sim$1040 cm-1 energetic splitting suggesting large electronic coupling between the sites. Note that unlike conformationally heterogeneous18, 34, 26 iSF dimers where Davydov splitting is inferred from FC-progression dominated linear spectra and DFT calculations, the Davydov splitting in the structurally well-defined rigid dimer investigated here is directly confirmed by the presence of 2DES cross-peaks.
@@
Figure 2: 2DES reveals Davydov splitting and excitation wavelength dependent $TT_{1}$ formation. (a) 2D spectrum of the NDI dimer at $T$ = 0.2 ps. The top panel shows the absorption spectrum (black) overlaid with the pump spectrum (gray shaded). The side panel shows the spectrally integrated spectrum versus the detection wavelength, $\lambda_{t}$, overlaid with the probe spectrum (gray shaded). Contours are drawn at 10% to 100% in 10% intervals. The positive ground state bleach (GSB) signal is shown in red while the negative excited state absorption (ESA) signal is shown in blue. (b) (left) Dynamics in the ESA region denoted by the green rectangle in panel A. (middle) Dynamics in the ESA region denoted by the yellow rectangle in panel A. (right) Dynamics in the ESA region denoted by dashed yellow rectangle in panel A. For panel b (left and middle), the data was collected until 5 ps and the spectrally integrated 2D signals are fitted to a freely floated 3-exponential model where the first two time-constants are shown in the figure. For panel b (right), the time constants of the fit function for the middle panel are kept fixed and only the amplitudes are floated. The accompanying top panel shows the respective residuals. The $\pm{\sigma}$ error band on the fit is calculated by averaging N=6 trials of $T$ = 0.5 ps 2D spectrum where collection of individual trials was interleaved over the duration of the full scan. The arrows denote the 200 fs decay or rise, that is followed by the slower rise of the ESA signal. (c) 2D rate maps obtained by global fitting of all 2D pixels to a 3-exponential model. The obtained time constants are shown above the respective maps. The rate maps are normalized by the sum of the maxima of two rate maps. Global rate analysis neglects the excitation wavelength dependence seen in panel b (left versus middle). Contours are drawn from 10% to 100% in 10% intervals.
@@
Interestingly, the strongest features in the 2D spectrum are the two distinct ESA bands along the excitation axis, where both approximately coincide with the position of the two Davydov components. As shown in Figure 2b(left, middle), when an area at these locations is sampled, green versus yellow rectangles in Figure 2a, a strong wavelength-dependent dynamics of the ESA signal is seen. The ESA signal shows a fast decay of 0.19 $\pm$ 0.05 ps and 0.17 $\pm$ 0.05 ps with green versus yellow excitation, which is approximately wavelength-independent within the error bars. The rise of the ESA signal is, however, $\sim$3x faster when the lower Davydov component is excited (2.12 $\pm$ 0.32 ps versus 0.60 $\pm$ 0.18 ps). The fitting procedure is described in Section S2 and summarized in Table S1, S2. The fits shown in Figure 2b (left, middle) are freely floated and do not show any trends in the residual (top panels in Figure 2b). In comparison, when the fit is constrained to have the same time constants, a clear residual trend is seen in the 650 nm excitation wavelength region. This is shown in Figure S3b(right). Figure 2b(right) shows the $T$ dynamics in the dashed yellow rectangle with excitation wavelength at the lower Davydov component as before, but with detection wavelength $\sim$710 nm compared to 760 nm earlier. The $\sim$200 fs decay changes to a $\sim$200 fs rise of the signal, followed by a slower rise on the $\sim$2 ps timescale. Previous32 low temporal resolution PP experiments identified the picosecond rise of the ESA signal as the signature of $TT_{1}$ formation. Here, we find that this process is strongly wavelength dependent, along with an intermediate step that occurs on a $\sim$200 fs timescale, which, however, is not wavelength dependent. Note that we have confirmed that all the above 2DES observations remain unchanged even if the experiments were conducted in a non-polar (weakly polar) DCM solvent ( Section S5) suggesting that the intermediate states during $S_{1}\rightarrow TT_{1}$ do not involve any significant $CT$ character (vide infra).